# DHA: Learning Decoupled-Head Attention from Transformer Checkpoints via Adaptive Heads Fusion

### v1

@@ 1 Introduction
With the limited understanding of parameter characteristics in modern LLMs, we first perform an empirical analysis from the perspectives of heads’ representation similarity. We observe that there are some head-clusters with high internal similarity in MHA checkpoints. Similar head clusters imply a enormous redundancy in MHA, which coincides with the sparsity found in previous studies [12, 13]. In particular, the clusters of key heads and value heads across different layers show a decoupled distribution, meaning that there is a significant variation in the distribution of head-cluster similarities across layers, key heads and value heads, as illustrated in Fig. 2(a),2(b). Intuitively, we can prune redundant heads based on the above characteristics. Nonetheless, each head has its unique role, and thus no heads should be arbitrarily discarded. Furthermore, we find that linear fusion based on multiple similar heads can reconstruct the original head functionality without causing a significant performance drop (see Sec. 3.1). Based on this observation, we believe that selectively fusing corresponding heads in clusters can construct a more efficient architecture with minimum loss.
@@
Figure 2: Visualization of the similarity between heads within the MHA of LLaMA2-7B model at the 0th layer (a) and the 21st layer (b). Details in Appendix D.1. Key heads and value heads exhibit decoupled distributions.
@@ 5.1 Experimental Setup
Our experimental framework utilizes the Sheared-LLaMA codebase [16] implemented on the Composer package [20], and is executed on 8 NVIDIA A100 GPUs (80GB). The models are trained with a sequence length of 4096, employing a global batch size of 64 during the fusion phase and 256 during the continued pre-training phases. DHA models were trained for 1000 steps (0.2B token budget) during the fusion phases For the continued pre-training, we trained both baseline models and DHA for up to 5000 steps (5B token budget). The learning rates were set at 1e-4 for language modeling loss, and 1e-2 for Lagrangian multipliers and fusion operators respectively.
@@ 5.3 Analysis
Allocating more computation to the fusion phase aids in better retention of information within the checkpoint. Our experiments assessed the effects of budget allocations between the fusion and CT phases within a fixed budget of 2 billion tokens. Tab. 3 shows that increasing the fusion budget consistently from 0.05B to 0.2B improves model performance at the initialization point. Training with just 0.1B data is sufficient to achieve a good starting point, and increasing fusion budget will not affect the final performance.
@@
For interpretability analysis, we visualized the parameter characteristics of the post-fusion DHA model in Fig. 9 (detials in Appendix D.2), and compared them with those prior to fusion. The DHA parameter distribution shows consistency with MHA’s. This indicates that DHA effectively aggregates multiple similar functional heads within clusters and new fused heads successfully reconstruct the functionalities of multiple origin heads in MHA. It is noteworthy that the significant reduction in the number of similar heads within the DHA architecture indicates that our method effectively reduces redundancy among the heads.
@@ Acknowledgments
We would like to thank Yinqi Yang, Jiawei Sheng, Xinhua Zhang, Shicheng Wang, Chuanyu Tang and members of the IIE KDsec NLP group for their valuable feedback and discussions. We are very grateful to Mengzhou Xia for providing the concise and effective ShearingLLaMA experimental code and for her assistance during the reproduction process. This research is supported by the National Key Research and Development Program of China (grant No.2021YFB3100600) and the Youth Innovation Promotion Association of CAS (Grant No. 2021153).
@@ A.1 Extended Related Works
KV Cache Compression methods emerged for reducing the prominent inference bottleneck caused by KV cache, particularly for long content input. A series of methods [8, 10, 9, 47] explored the sparsity among Transformer’s attention block, then evicted unnecessary tokens from KV Cache for efficient inference. However, these methods discard information from the context and use algorithms for inference that are inconsistent with the training phase, which can cause model performance degradation. DHA does not need to discard information from the context and is able to maintain consistent performance for training and inference. Pruning [15, 16], quantization [48, 49, 50] and distillation [51, 52, 53] can reduce the number of model key and value headers, parameter dimensions, and activation to reduce memory bandwidth overhead during model inference. Deja Vu [54] and CHAI [7] prune pruning redundant heads through clustering methods for efficient inference. In the LLM era, this leads to a significant reduction in neuron redundancy as models move from task-specific to generalized [55]. The application of these methods to LLMs is computationally expensive and leads to performance degradation at larger pruning magnitudes.
@@ B.4 Layer Allocation Based on Fusion Loss
This algorithm efficiently allocates resources to different layers based on their respective losses to optimize system performance. Initially, it assigns a minimum allocation to each layer. Then, it calculates weights for each layer based on their losses, prioritizing layers with higher losses. The algorithm determines the number of times 16 can be allocated based on the remaining allocation. It allocates 16s to layers with the highest weights until reaching a predetermined limit. Next, it redistributes the remaining allocation to layers with the highest loss-to-allocation ratios, assigning resources in multiples of 8 or 4. This process ensures that layers with higher losses receive more resources, optimizing the overall system performance. Finally, the algorithm returns the final allocation for each layer, resulting in an efficient distribution of resources across the system.
@@ B.5 Training Details
The hyperparameters used in our experiments are presented in Table 4. We employ fully sharded data parallel to efficiently train our models in parallel, and we utilize FlashAttention V1 [58] to accelerate the training process. A cosine learning rate scheduler is used, with the learning rate decaying to a minimum of 10% of the peak value. Preliminary experiments were conducted to determine the optimal peak learning rate for learning the fusion variables and Lagrange multipliers.
@@ Appendix C Extended Experiments
As shown in Figure 10, the tuned DHA model outperforms all GQA baselines of comparable scale . This demonstrates that the DHA model effectively retains the foundational capabilities of the MHA model and can be activated through instruction tuning to produce long, coherent, and informative responses.
@@
Appendix D Extend Observation
D.1 Header parameter characteristics in MHA
@@
D.2 Header parameter characteristics in DHA

### v2

@@ 1 Introduction
With the limited understanding of parameter characteristics in modern LLMs, we first perform an empirical analysis from the perspectives of heads’ parameter similarity. We observe that there are some head-clusters with high internal similarity in MHA checkpoints. Similar head clusters imply a enormous redundancy in MHA, which coincides with the sparsity found in previous studies [12, 13]. In particular, the clusters of key heads and value heads across different layers show a decoupled distribution, meaning that there is a significant variation in the distribution of head-cluster similarities across layers, key heads and value heads, as illustrated in Fig. 2(a),2(b). Intuitively, we can prune redundant heads based on the above characteristics. Nonetheless, each head has its unique role, and thus no heads should be arbitrarily discarded. Furthermore, we find that linear fusion based on multiple similar heads can reconstruct the original head functionality without causing a significant performance drop (see Sec. 3.1). Based on this observation, we believe that selectively fusing corresponding heads in clusters can construct a more efficient architecture with minimum loss.
@@
Figure 2: Visualization of the similarity between heads within the MHA of LLaMA2-7B model at the 0th layer (a) and the 21st layer (b). Details in Appendix E.1. Key heads and value heads exhibit decoupled distributions.
@@ 5.1 Experimental Setup
Our experimental framework utilizes the Sheared-LLaMA codebase [16] implemented on the Composer package [20], and is executed on 8 NVIDIA A100 GPUs (80GB). The models are trained with a sequence length of 4096, employing a global batch size of 64 during the fusion phase and 256 during the continued pre-training phases. The learning rates were set at 1e-4 for language modeling loss, and 1e-2 for Lagrangian multipliers and fusion operators respectively.
Budget.
DHA models were trained for 1000 steps (0.2B token budget) during the fusion phases. For the continued pre-training, we trained both baseline models and DHA for up to 50000 steps (50B token budget). To evaluate the training acceleration capability of DHA, we evaluate its performance under two budget scenarios. First, we set a budget of 1B tokens to compare the early-stage rapid convergence capabilities of DHA and GQA. Then, we set a budget of 50B tokens to further assess the performance of DHA over a more extended training period.
@@ 5.3 Analysis
Allocating more computation to the fusion phase aids in better retention of information within the checkpoint. Our experiments assessed the effects of budget allocations between the fusion and CT phases within a fixed budget of 2 billion tokens. Tab. 3 shows that increasing the fusion budget consistently from 0.05B to 0.2B improves model performance at the initialization point. Training with just 0.1B data is sufficient to achieve a good starting point, and increasing fusion budget will not affect the final performance. This experiment also demonstrates the necessity and effectiveness of the fusion stage under low-resource conditions. When we have a larger training budget, we can allocate more resources to the fusion stage to achieve a better initialization point for DHA.
@@
For interpretability analysis, we visualized the parameter characteristics of the post-fusion DHA model in Fig. 9 (detials in Appendix E.2), and compared them with those prior to fusion. The DHA parameter distribution shows consistency with MHA’s. This indicates that DHA effectively aggregates multiple similar functional heads within clusters and new fused heads successfully reconstruct the functionalities of multiple origin heads in MHA. It is noteworthy that the significant reduction in the number of similar heads within the DHA architecture indicates that our method effectively reduces redundancy among the heads.
@@ Acknowledgments
We would like to thank Yinqi Yang, Jiawei Sheng, Xinhua Zhang, Shicheng Wang, Chuanyu Tang and members of the IIE KDsec NLP group for their valuable feedback and discussions. We are very grateful to Mengzhou Xia for providing the concise and effective ShearingLLaMA experimental code and for her assistance during the reproduction process. Work done during Yilong Chen’s internship in Baidu Inc. This research is supported by the National Key Research and Development Program of China (grant No.2021YFB3100600) and the Youth Innovation Promotion Association of CAS (Grant No. 2021153).
@@ A.1 Extended Related Works
KV Cache Compression methods emerged for reducing the prominent inference bottleneck caused by KV cache, particularly for long content input. A series of methods [8, 10, 9, 47, 48] explored the sparsity among Transformer’s attention block, then evicted unnecessary tokens from KV Cache for efficient inference. However, these methods discard information from the context and use algorithms for inference that are inconsistent with the training phase, which can cause model performance degradation. DHA does not need to discard information from the context and is able to maintain consistent performance for training and inference. Pruning [15, 16], quantization [49, 50, 51] and distillation [52, 53, 54] can reduce the number of model key and value headers, parameter dimensions, and activation to reduce memory bandwidth overhead during model inference. Deja Vu [55] and CHAI [7] prune pruning redundant heads through clustering methods for efficient inference. In the LLM era, this leads to a significant reduction in neuron redundancy as models move from task-specific to generalized [56]. The application of these methods to LLMs is computationally expensive and leads to performance degradation at larger pruning magnitudes.
@@ B.4 Layer Allocation Based on Fusion Loss
This algorithm efficiently allocates resources to different layers based on their respective losses to optimize system performance. Initially, it assigns a minimum allocation to each layer. Then, it calculates weights for each layer based on their losses, prioritizing layers with higher losses. The algorithm determines the number of times 16 can be allocated based on the remaining allocation. It allocates 16s to layers with the highest weights until reaching a predetermined limit. Next, it redistributes the remaining allocation to layers with the highest loss-to-allocation ratios, assigning resources in multiples of 8 or 4. This process ensures that layers with higher losses receive more resources, optimizing the overall system performance. Finally, the algorithm returns the final allocation for each layer, resulting in an efficient distribution of resources across the system. The total search process for the LLaMA2 model requires 42 minutes.
@@ B.5 Training Details
The hyperparameters used in our experiments are presented in Tab. 4. We employ fully sharded data parallel to efficiently train our models in parallel, and we utilize FlashAttention V1 [59] to accelerate the training process. A cosine learning rate scheduler is used, with the learning rate decaying to a minimum of 10% of the peak value. Preliminary experiments were conducted to determine the optimal peak learning rate for learning the fusion variables and Lagrange multipliers.
@@ Appendix C Extended Experiments
As shown in Fig. 10, the tuned DHA model outperforms all GQA baselines of comparable scale . This demonstrates that the DHA model effectively retains the foundational capabilities of the MHA model and can be activated through instruction tuning to produce long, coherent, and informative responses.
@@
Combination with KV Cache Compression Techniques.
In Sections 2 and 4, we demonstrated that DHA is a more efficient GQA architecture, so it has similarly good compatibility. We tested the compatibility of the DHA model with the KVCache eviction method NACL [48]. NACL 25% indicates retaining only 25% of the KVCache. The experiment results are shown in the Tab. 5. DHA and GQA exhibit equally good compatibility with KV cache compression techniques.
Table 5: Comparison of log(PPL) between DHA and GQA with NACL.
Method log(PPL)
GQA-7b-25% 2.89
DHA-7b-25% 2.84
GQA-7b-25% (NACL 25%) 3.01
DHA-7b-25% (NACL 25%) 2.93
Table 6: Comparison of Avg ACC and PPL between different methods at 7B-25% (5B).
Method Avg ACC PPL
DHA-7B-25% (5B) 62.4 7.29
GQA-7B-25% (5B) 60.3 7.54
GQA (CKA-Grouping)-7B-25% (5B) 60.4 7.51
Compare with Advance GQA Initialization.
It’s a common and effective approach to convert MHA to GQA using mean pooling instead of training from scratch. The author of GQA tested several methods for the initialization of GQA and found it works best using simple mean pooling from MHA. Indeed, training GQA from scratch will cost trillions tokens budget to match the performance of MHA which is inefficient and costly.Inspired by the similarity of head parameters, we improved the initialization method of GQA: instead of direct grouping, we first cluster similar heads using CKA and then perform mean-pooling initialization within each cluster. We compare this approach with the Vanilla GQA and DHA.
Tab. 6 shows that GQA(CKA-Grouping)-7B-25% (5B) achieved comparable performance to the original implementation in Vanilla GQA. We believe the reason for this is that the head grouping learned by DHA is based on the fusible nature between heads, which cannot be completely equated with CKA similarity. More importantly, DHA not only groups heads based on similarity but also learns the fusible parameters. This allows it to eliminate the influence of redundant parameters and retain more important information during the initialization process, which is not possible with mean initialization.
Appendix D Extend Analysis
How Merging Weights Change.
Refer to Fig. 3(a), where we show the weight variation diagram. In the fusion process of heads 0-3, head 0 initially constitutes 100% as the starting head of the MHA. As the fusion process progresses, the parameters of the important heads increase, and the proportions of all heads become more balanced. This indicates that the algorithm attempts to retain information from different heads by balancing the parameter proportions of each head. This process results in a slight increase in loss, but not significantly.
DHA’s Compatibility on GQA Model.
DHA is primarily designed for models based on the Transformer Decoder architecture and can be adapted to all models with this architecture. We chose LLaMA [60] as the experimental baseline because it is a classic model using the decoder architecture in LLMs. Other open-source LLM models differ from LLaMA only in certain details (such as activation functions and training methods), which do not affect DHA’s training. Successfully applying DHA to LLaMA indicates that it can be used in most decoder-only models. GQA [5] is an efficient variant of MHA, which optimizes the inference process through head grouping and sharing. Due to its simplicity and efficiency, GQA is widely used. DHA can be similarly constructed based on GQA, requiring only minor adjustments to the construction process. Here, we provide two feasible methods to convert GQA to DHA.
  • Easiest method in less than 1 minute. GQA can be losslessly converted into MHA by simply replicating the GQA’ KV heads. Then, we can perform the DHA transformation on the MHA architecture.
  • Minor modification by grouping KV. DHA only needs to group and fuse the Key and Value heads. When constructing DHA on GQA, we initially group the Key and Value, maintaining alignment with GQA functionality. During the training phase, the fused head parameters can replace the original GQA heads for sharing.
Inter-layer Grouping of Heads or Only Intra-layer Grouping?
Only intra-layer grouping and fusion is conducted in DHA. Fig. 1 meant to illustrate the decoupled-heads where the number of key and value heads can be different among layers. The DHA method employs parameter fusion within each layer for three reasons:
  • Higher redundancy of heads within layer for fusion. The heads within a layer exhibit high similarity and redundancy, which provides a good starting point for parameter fusion.
  • More complex optimization for inter-layer fusion. The optimization process between layers is very complex and requires memory operations for cross-layer calls, which inherently increases the inference cost.
  • Promising future work by introducing inter-layer fusion [61]. This paper represents an early exploration of applying parameter fusion methods within model parameters. The inter-layer fusion approach is indeed a valuable direction for future exploration.
Accuracy Loss after Transformation.
The performance gap between the results shown in the paper and MHA is primarily due to the following two reasons:
  • The gap of pre-training data. The MHA model was not trained on the same data used for DHA. Since LLaMA’s training data is not directly open-sourced, we used an experimental open-sourced pre-training data following Sheared-LLaMA (Xia et al., 2024). The improved pre-training data will close the gap between DHA and MHA.
  • Parameter size difference. Compared to MHA, DHA compresses 50% or 25% of attention heads, requires only 0.05% of pre-training data and achieves approximately 5% loss. The number of parameters of MHA is much larger than that of DHA, so performance loss is inevitable during conversion. Compared with GQA, a strong baseline with the same number of parameters, DHA has shown higher training efficiency and performance advantages. Due to the high efficiency of DHA, DHA can use more heads than MHA with the same number of parameters, and has the opportunity to achieve better performance.
Appendix E Extend Observation
E.1 Header parameter characteristics in MHA
@@
E.2 Header parameter characteristics in DHA
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract provides a concise summary of the key findings and experiment results. The introduction in Sec. 1 outlines the research questions and objectives in paragraph 3,4 and contribution in paragraph 5.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper discusses the limitations of the work performed by the authors in detail in Appendix Appendix. A.2, highlighting two specific limitations and the broader impact.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: This paper is mainly based on observation, making conjectures and methods and proving the effects through experiments. The paper defines the background in Sec. 2, presents the conjecture in Sec. 3, and provides a detailed derivation of the form and optimization process in Sec. 4. All assumptions made in the paper are thoroughly validated through experiments in Sec. 5.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper provides a detailed description of the experimental data setup and hyperparameter settings in Sec. 5. Additionally, in Sec. 4 and in Appendix. B.2 sections we thoroughly explain the derivation and implementation process, ensuring all necessary information for reproducing the main experimental results is disclosed.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: The datasets, baseline methods, and models used in the paper are fully open-source and available on Hugging Face. The paper includes the key implementation steps and code in Sec. 4 and the Appendix. B.2. However, the complete code is still being organized and is under consideration for open sourcing.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes] ,
  29. Justification: The paper provides a detailed description of the experimental data setup and hyperparameter settings in Sec. 5. Additionally, in Sec. 4 and in Appendix. B.2 sections we thoroughly explain the derivation and implementation process.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All results are averaged over multiple tests, and we report the mean accuracy along with the standard deviation (acc_norm) as a measure of error bars.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In Sec. 5.1, we report the GPUs we used, the memory, and detailed training information. For more information you can refer to the Appendix B.2.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer:[Yes]
  44. Justification: The discussion of the ethics and impact can be consulted in Appendix. A.2. We are open and transparent throughout the study and do not design for human subjects, privacy data bias, or other issues.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The discussion of the broader impacts can be consulted in Appendix. A.2.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper presents an improved approach based on the existing model architecture, but does not release any new models. The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: This article uses assets reasonably in compliance with the license, and the assets used are cited in the article.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.